OpenStudy (anonymous):

A long distance trucker traveled 224 miles in one direction during a snow storm. The return trip in rainy weather was accomplished at double the speed and took 4 hour less time. Find the speed going. The speed going was ? mph.

OpenStudy (anonymous):

remember distance = speed times time, so time T = distance D/speed S time outbound T = 224/S inbound T-4 = 224/2S Solve simultaneous equations by substitution to get S

OpenStudy (amistre64):

i dint do it right lol

OpenStudy (amistre64):

st = 224; t = 224/s 2s(t-4) = 224 2st -2s4 = 224 2s(224) ------- - 8s = 224 s 448 - 8s = 224 224 = 8s..... t=8 s = 28 going

OpenStudy (amistre64):

:) when you double the speed; you half the time in the same distance...